# From Flatland to Space: Teaching Vision-Language Models to Perceive and Reason in 3D

### v4

@@ 2 Related works
Existing multimodal datasets primarily provide single-view or limited multi-view images, often designed for image-text retrieval or basic visual question answering, making them insufficient for advanced 3D reasoning [21, 22, 23, 24]. Some datasets attempt to address this by collecting large-scale images from the internet, but they lack precise 3D annotations, relying instead on estimation models, which limits their effectiveness in multi-view and spatial reasoning tasks [6, 7]. Others incorporate image sequences or 3D scene representations for object counting, spatial relationship reasoning, or grounding, but they mainly focus on high-level semantics while neglecting low-level geometric cues such as depth, distance, and camera parameters [13, 14, 25]. Recent efforts [26] explore spatial reasoning from a robotics perspective, emphasizing reference-frame understanding and grounded spatial relations in real-world scenes. While highly relevant to embodied agents, these datasets focus on a narrower set of reasoning types and lack the task diversity and cognitive stratification we aim for. To overcome these limitations, we propose a 3D-ground-truth-based data generation pipeline capable of rendering single-view, multi-view, and video-based data. Our approach ensures precise spatial supervision, covering a broad spectrum of tasks from basic perception to advanced reasoning, structured into three difficulty levels. This provides a more comprehensive and scalable solution compared to existing datasets.

### v5

@@ 2 Related works
Existing multimodal datasets primarily provide single-view or limited multi-view images, often designed for image-text retrieval or basic visual question answering, making them insufficient for advanced 3D reasoning [21, 22, 23, 24]. Some datasets attempt to address this by collecting large-scale images from the internet, but they lack precise 3D annotations, relying instead on estimation models, which limits their effectiveness in multi-view and spatial reasoning tasks [6, 7]. Others incorporate image sequences or 3D scene representations for object counting, spatial relationship reasoning, or grounding, but they mainly focus on high-level semantics while neglecting low-level geometric cues such as depth, distance, and camera parameters [13, 14, 25]. Concurrently, datasets like MSR3D [26] are also emerging, offering large-scale, multi-modal inputs that explicitly include point clouds alongside text and images. These datasets are designed to evaluate situated reasoning and navigation within 3D environments, providing valuable benchmarks for models that directly process 3D data. Recent efforts [27] explore spatial reasoning from a robotics perspective, emphasizing reference-frame understanding and grounded spatial relations in real-world scenes. While these approaches, along with those directly utilizing 3D inputs like MSR3D, are invaluable for specific applications such as embodied AI and situated interaction, our work takes a different stance. We aim to achieve a broader spectrum of general 3D spatial tasks and a cognitive stratification of reasoning abilities, learned primarily from 2D visual data derived from 3D ground truth. This explores the extent to which VLMs can develop rich 3D understanding without explicit 3D input modalities during inference. To overcome these limitations and to further explore this potential, we propose a 3D-ground-truth-based data generation pipeline capable of rendering single-view, multi-view, and video-based data. Our approach ensures precise spatial supervision, covering a broad spectrum of tasks from basic perception to advanced reasoning, structured into three difficulty levels. This provides a more comprehensive and scalable solution that complements approaches leveraging direct 3D representations or focusing on specific robotic interactions.